OpenStudy (anonymous):

How does the square root of 4 + 16 turn into 2 times the square root of 5? www.mathportal.org/calculators/analytic-geometry/distance-and-midpoint-calculator.php?val1=3&val2=-2&val3=1&val4=-6&rb1=dist ( at the bottom )

OpenStudy (michele_laino):

please note that: \[\sqrt{4+16}=\sqrt{20}=\sqrt{4*5}=\sqrt{2^{2}*5}=2\sqrt{5}\]

OpenStudy (michele_laino):

@mizuno

OpenStudy (anonymous):

Thanks! @Michele_Laino

OpenStudy (michele_laino):

Thanks! @mizuno